# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@ 3.5 Labelling the shocks
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroskedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017). Additionally, in line with the statistical identification literature, any economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
@@
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, because the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, due to the problem of weak identification with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, discussed in Sections 3.3 and 3.4, the ordering and signs of the columns of $B_{2},...,B_{M}$ is often practically arbitrary even when the ordering and signs are, in theory, uniquely identified by Proposition 1 or 2.
Assuming that the parameters are estimated by the method of maximum likelihood (ML), weak identification with respect to the ordering and signs of the columns often results in multiple local maxima of the log-likelihood function with log-likelihoods very close to each other. Our estimation method, discussed in Section 4 and implemented to the R package sstvars (Virolainen, 2024), can produce a set of local solutions to the estimation problem. This allows the researcher to check whether there are multiple local solutions with log-likelihoods very close to each other. In practice, one can then simply choose from the almost equally well fitting local solutions the one that, based on economic reasoning, facilitates labelling the shock(s) of interest to the same column of the impact matrix in all regimes.
To exemplify, consider the simplistic setup of two variables, an output variable and a price variable, two shocks, the aggregate supply shock and the aggregate demand shock, and two regimes ($M=2$). Suppose the estimates of the impact matrices of the two regimes with the highest log-likelihood are the following:
$\hat{B}_{1}=\begin{bmatrix}0.67&\phantom{-}0.21\\ 0.06&-0.21\\ \end{bmatrix}\ \ \text{and}\ \ \hat{B}_{2}=\begin{bmatrix}-0.55&1.14\\ \phantom{-}0.49&0.09\\ \end{bmatrix}.$ (3.5)
Then, suppose there exists another local solution solution with the second highest likelihood that is almost equal to the highest log-likelihood, but the estimates of the impact matrices are the following:
$\hat{B}_{1}=\begin{bmatrix}0.59&\phantom{-}0.16\\ 0.05&-0.17\\ \end{bmatrix}\ \ \text{and}\ \ \hat{B}_{2}=\begin{bmatrix}1.05&\phantom{-}0.58\\ 0.18&-0.41\\ \end{bmatrix}.$ (3.6)
In the estimates (3.5), the first shock moves output and prices in the same direction at impact in Regime 1 but in opposite directions in Regime 2. The second shock in, turn, moves output and prices in opposite directions in Regime 1 but in the same direction in Regime 2. Thus, neither of the shocks can be labelled as the aggregate supply shock nor as the aggregate demand shock. On the other hand, in the estimates (3.6), the first shock moves output and prices in the same direction satisfies the same signs of the impact effects in both regimes, so it can be labelled as the aggregate supply shock. The second shock moves output and prices in opposite directions and satisfies the same signs of the impact effects in both regimes, so it can be labelled as the supply shock.
Since the estimates in (3.5) do not align with what is expected from the aggregate supply and demand shocks but the estimates in (3.6) do, the latter solution can be chosen based on economic reasoning despite having marginally smaller log-likelihood. [^1] It is also possible to obtain the estimates in (3.6) more formally by using restricted ML estimation, imposing the restriction that the first shock moves both variables in the same direction and the second shock in opposite directions on impact. Since the solution (3.6) yields the second highest log-likelihood among all local solutions, it is obtained from the restricted ML estimation. Note that also other types of information than just signs of the impact responses of the variable can be utilized to label the shocks. For instance, if the shock of interest should or should not significantly move certain variables on impact, the corresponding estimates of $B_{1},...,B_{M}$ should satisfy this property. Lanne et al. (2017) go further and exploit more information by labelling the shocks based on the estimated impulse response functions.
If none of the alternative local solutions with almost equal log-likelihoods allows the shock of interest to be labelled to same column across $B_{1},...,B_{M}$, the identification can be strengthened by imposing economic overidentifying restrictions on the model. For example, if economic reasoning suggests that certain variables should respond to the shock of interest with a lag, the appropriate zero restrictions can be imposed on the corresponding entries of $B_{y,t}$. Due to the statistical identification, these zero restrictions are overidentifying and thereby testable, which a major advantage compared to using conventional short-run restrictions alone.
@@ 4 Estimation
The parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of maximum likelihood. Obtaining the log-likelihood function requires specifying a distribution for the structural errors $e_{t}=(e_{1t},...,e_{dt})$ satisfying Assumption 1. In our empirical application presented in Section 5, we assume that each shock $e_{it}$, $i=1,...,d$, follows a Student’s $t$ distribution with mean zero, variance one, and degrees of freedom $\nu_{i}\in(2,\infty)$. The assumption $\nu_{i}>2$ is made to ensure the existence of the second moments, and the assumption $\nu_{i}<\infty$ is made to distinguish the distribution from the Gaussian distribution.
In order to normalize the variances of the structural shocks to unity, we parametrize the Student’s $t$ distribution using its variance as a parameter together with the mean and degrees-of-freedom. Similar parametrization has been utilized previously in Virolainen (2022) and Meitz et al. (2023), for example. The density function of such a univariate $t$-distribution with mean $\mu_{i}$, variance $\sigma_{i}^{2}>0$, and $\nu_{i}>2$ degrees-of-freedom is
$t_{1}(y_{it};\mu_{i},\sigma_{i}^{2},\nu_{i})=\frac{\Gamma\left(\frac{1+\nu_{i}}{2}\right)}{\sqrt{\pi(\nu_{i}-2)}\Gamma\left(\frac{\nu_{i}}{2}\right)}\sigma_{i}^{-1}\left(1+\frac{(y_{it}-\mu_{i})^{2}}{(\nu_{i}-2)\sigma_{i}^{2}}\right)^{-(1+\nu_{i})/2},$ (4.1)
where $\Gamma\left(\cdot\right)$ is the gamma function.
We collect the parameters to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{m,0},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),....,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(vec(B_{1}),...,vec(B_{M}))$, $\alpha$ contains the transition weight parameters (and is omitted if exogenous weights are used), and $\nu=(\nu_{1},...,\nu_{d})$ contains the degrees-of-freedom parameters. Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
$L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}t_{1}(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{t});0,1,\nu_{i})\right),$ (4.2)
where the impact matrix $B_{y,t}$ is defined in Equation (3.3), $I_{d,i}$ is the $i$th column of the $(d\times d)$ identity matrix $I_{d}$, and the Student’s $t$ density $t_{1}(\cdot;0,1,\nu_{i})$ is defined in (4.1). The ML estimator maximizes the log-likelihood function (4.2) with respect to the parameter $\boldsymbol{\theta}$.
@@
To facilitate obtaining desirable asymptotic properties for the ML estimator, such as consistency and a conventional limiting distribution, it is useful to establish the ergodicity and stationarity of the structural STVAR model. Based on the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one. However, since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in a numerical ML estimation. Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition. After the estimation, the sufficient condition can be checked for the solutions of interest.
In practice, maximizing the log-likelihood function typically requires numerical optimization techniques. Particularly if the transition weights are endogenous, there are often a very large number of modes in the surface of the log-likelihood function, and large areas in the parameter space where it is flat in multiple directions. To address these complexities, we follow Meitz et al. (2023) and Virolainen (2022, forthcoming) and adopt a two-phase estimation procedure, which is ran a large number of times. In the first phase, a genetic algorithm is employed to find starting values for a gradient based variable metric algorithm. In the second phase, the variable metric algorithm is then initialized from each starting value found in the first phase, thereby often converging to a nearby local maximum. Some of the estimation rounds may end up in saddle points or points near the boundary of the parameters space (e.g., in points where $B_{y,t}$ is nearly singular for some $t$, or Condition A.1 in Appendix A, if imposed, is close to breaking), and some of the local solutions might be inappropriate for statistical inference (e.g., there might be only few observations from some of the regimes).
As the log-likelihood function is highly multimodal and the estimation algorithm is ran a large number of times, it often results in a variety of local solutions. After the estimation rounds have been ran, the solution that maximizes the log-likelihood function among the appropriate local solutions can be selected. Then, if the transition weights are endogenous and a stationary solution is preferred, the sufficient condition for ergodic stationarity of the model (Assumption A.1(I) in Appendix A) can be checked for the local solutions of interest to select the best solution whose stationarity can be verified. If there are multiple such solutions with almost equal log-likelihoods, the best local solution that facilitates labelling the shocks of interest can selected based on the procedure discussed in Section 3.5. The accompanying CRAN distributed R package sstvars (Virolainen, 2024) implements the two-phase estimation procedure utilizing a modified genetic algorithm.
@@ 5 Empirical application
Our empirical application studies the macroeconomic effects of the climate policy uncertainty shock. Currently, there exist only a few studies on the subject. Fried et al. (2021) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy risk decreases total capital stock, causing output and carbon emissions to decrease. Gavriilidis et al. (2023), in turn, propose a news based index on measuring climate policy uncertainty (CPU), and the study the effects of the CPU shock using a linear recursively identified SVAR model and local projections. They find that a positive shock to their CPU index (CPUI) decreases real activity and increases prices, thus, transmitting to the U.S. economy as a supply shock.
We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty. Such variation can be accommodated by our structural STVAR model by specifying a logistic transition weight function with the lagged level of economic policy uncertainty as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more nuanced results on the macroeconomic effects of the CPU shock than linear SVAR analysis.

### v3

@@ 3.5 Labelling the shocks and blended identification
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017). Additionally, in line with the statistical identification literature, many economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
@@
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, as discussed in Sections 3.3 and 3.4, the identification is often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other. If the identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$. As our estimation method, discussed in Section 4, can produce a set of local solutions, the possible presence of such local solutions can be easily checked.
To formally obtain the local solution that facilitates labelling the shock of interest to the same column of the impact matrix in all regimes, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with additional identifying information (cf. Carriero et al., 2024). As a concrete example, if a supply shock should increase output and decrease prices at impact in all regimes, selecting the best local solution where the same column in all $B_{1},...,B_{M}$ satisfies such signs amounts to combining identification by non-Gaussianity with the appropriate short-run sign restrictions. In addition to short-run sign restrictions, also other types of information can be exploited in blended identification, including narrative restrictions and testable zero impact effect restrictions, for instance.
@@ 4 Estimation
4.1 Maximum likelihood estimation and ergodic stationarity
@@
The parameters of the structural STVAR models discussed in Section 3 can be estimated by the method of maximum likelihood (ML), which amounts to maximizing the log-likelihood function over the parameter space. To obtain a well-behaving estimator with desirable asymptotic properties such as consistency, the parameter space is often restricted to the region where the model is ergodic stationary. Based on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one. But since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in estimation. Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition. After estimation, the sufficient condition can be checked for the solutions of interest.
Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics. Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum. To overcome this issue, we propose allowing for unstable estimates, but adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close the boundary of the stability region, and then maximizing the penalized log-likelihood function. In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance of the optimization algorithms, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
Penalized likelihood -based estimation has been previously applied to time series models by Nielsen and Rahbek (2024). They impose nonnegativity constraints in the autoregressive conditional heteroscedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space. Nielsen and Rahbek (2024) use penalization also to determine which parameters should be set to zero.
4.2 The penalized log-likelihood function
Before introducing the penalized log-likelihood function, the log-likelihood function is presented. We assume that each shock $e_{it}$, $i=1,...,d$, follow the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$. The density function of the skewed $t$-distribution, $st(\cdot;\nu_{i},\lambda_{i})$, is given as (Hansen, 1994, Equations (10)-(13)):
$st(e_{it};\nu_{i},\lambda_{i})=b_{i}c_{i}\left(1+\frac{1}{\nu_{i}-2}\left(\frac{b_{i}e_{it}+a_{i}}{1-\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}\lambda_{i}+\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}\lambda_{i}}\right)^{2}\right)^{-(\nu_{i}+1)/2}$ (4.1)
where $\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}<-a_{i}/b_{i}$ and zero otherwise, and $\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}\geq-a_{i}/b_{i}$ and zero otherwise. The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is Gamma function.
Indexing the observed data as $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, the conditional log-likelihood function, conditional on the initial values $\boldsymbol{y}_{0}=(y_{0},...,y_{-p+1})$, is given as
$L(\boldsymbol{\theta})=\sum_{t=1}^{T}\log\left(|\det(B_{y,t})|^{-1}\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{y,t}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})\right),$ (4.2)
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3). The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{M,0},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degrees-of-freedom and skewness parameters. [^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights $\alpha$ is omitted from the parameter vector.
The penalized likelihood function is then defined as
$PL(\boldsymbol{\theta})=L(\boldsymbol{\theta})-P(\boldsymbol{\theta}),$ (4.3)
where $L(\boldsymbol{\theta})$ is defined in (4.2) and $P(\boldsymbol{\theta})\geq 0$ is a penalization term that penalizes the log-likelihood function from parameter values that are close to entering or are in an uninteresting region of the parameter space. Focusing on avoiding unstable estimates, we define the penalization term as
$P(\boldsymbol{\theta})=\kappa Td\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (4.4)
where $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of Regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization.
Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues. Our penalization term incorporates the multiplicative coefficient $Td$ in attempt to standardize the strength of the penalization with respect to the number of observations $T$ and variables $d$, as the value of the log-likelihood function (4.2) tends to scale with these quantities. In our empirical application (Section 5), we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, thus, penalizing parameter values outside the stability region significantly, while maintaining flexibility near the boundary.
4.3 Three-phase estimation procedure
Lanne et al. (2017) propose a three-phase estimation procedure for their linear SVAR model identified by non-Gaussianity. In the first phase, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent phases to maximize the log-likelihood function. The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function. To address these challenges, we propose a modified version of the three-phase estimation procedure in which the multimodality is explicitly taken into account.
Our three-phase estimation procedure operates with the following steps:
  1. 1.
    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1,0},...,\phi_{M,0},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) as described in Appendix C. Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
  2. 2.
    Estimate the error distribution parameters $\boldsymbol{\theta}_{2}\equiv(\sigma,\nu)$ by (penalized) ML with a genetic algorithm conditionally on the (penalized) NLS estimate $\hat{\boldsymbol{\theta}}_{1}^{NLS}$. Denote the obtained estimate of $\boldsymbol{\theta}_{2}$ as $\hat{\boldsymbol{\theta}}_{2}^{GA}$.
  3. 3.
    Estimate the full parameter vector $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$ by (penalized) ML by initializing a gradient based optimization algorithm from $(\hat{\boldsymbol{\theta}}_{1}^{NLS},\hat{\boldsymbol{\theta}}_{2}^{GA})$.
Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al. (2017) but with the nonlinearity of the model explicitly accounted for. Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-phase procedure. Step 2 is comparable to Step 2 of Lanne et al. (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able escape from local maxima. Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results. Since multiple estimation rounds are ran, the estimation procedure produces a set of estimates, possibly corresponding to different local maximums of the (penalized) log-likelihood function. If there are multiple solutions with (penalized) log-likelihoods almost equal to the highest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest. The proposed three-phase estimation procedure is implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-phase estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D. Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$. To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
According to the results (Tables D.2 and D.3 in Appendix D), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias. Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models. While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for a degrees-of-freedom parameter, and for a certain entry in each impact matrix, explainable by the statistical properties of the true data generating process. However, the bias diminishes for all of the parameters as the sample size increases. Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator. Overall, the performance of the proposed PML estimator and the three-phase estimation procedure seems reasonable.
@@ 5 Empirical application
Our empirical application studies the macroeconomic effects of the climate policy uncertainty (CPU) shock. In the related literature, Fried et al. (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output. Khalil and Strobel (2023) reach similar conclusions in their DSGE model, attributing the effects to financial institutions’ aversion to uncertainty. Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.